OpenStudy (anonymous):

Question: Estimate the limit: lim lnx/x-1 as x goes to 1

OpenStudy (anonymous):

do you know lhopitals rule?

OpenStudy (anonymous):

we talked about it earlier this week in class but I didn't get a clear understanding.

OpenStudy (anonymous):

this is LHopital's rule: \(\huge \lim_{x\rightarrow1} \frac{lnx}{x-1}=\lim_{x\rightarrow1} \frac{[lnx]'}{[x-1]'} \)

OpenStudy (anonymous):

ok but when he explained this it made no sense

OpenStudy (anonymous):

prolly because of the conditions needed in order to use the rule... to use this rule you have to determine if you have an indeterminate case: 0/0, inf/inf, 0^inf, etc.... this case is 0/0, so the rule applies...

OpenStudy (anonymous):

O ok

OpenStudy (anonymous):

so taking the derivative of the top: 1/x derivative of the bottom: 1 so lim(as x approaches 1) lnx/(x-1) = lim(as approaches 1) 1/x/1 = 1

OpenStudy (anonymous):

ok